# TIPO: Text to Image with Text Presampling for Prompt Optimization

### v2

@@ A.1 Danbooru2023
The Danbooru2023 dataset [65, 64, 45] is an extensive collection of images and their corresponding tags, compiled from the Danbooru image board. This dataset includes images annotated with highly specific and detailed tags, providing a rich resource for training both the Text-to-Image (T2I) and Large Language Models (LLMs) involved in the TIPO framework. The dataset contains data up to image ID 7,349,999, encompassing a wide variety of visual content with granular annotations. These annotations allow for the creation of nuanced and precise prompts, ensuring that longer, more detailed prompts can indicate subsets of shorter prompts.
@@
  • Rich Annotations: Danbooru2023’s detailed tags enable the differentiation of subtle variations in image content, such as “long hair” versus “very long hair,” which is crucial for the specificity required in high-quality image generation.
  • Large Volume: The dataset’s extensive size provides a robust foundation for training models, ensuring that they can learn from a diverse array of images and annotations.
  • Tag-Based Prompting: Utilizing the detailed tags from Danbooru2023, the LLM can generate refined prompts that lead to more accurate and high-quality image generation by the T2I model.
@@ A.2 GBC10M
The GBC10M dataset [29] is a large-scale collection of 10 million images sourced from CC12M [13], annotated using the Graph-Based Captioning (GBC) approach. Each image in GBC10M is represented by a graph where nodes correspond to object regions, compositions, and relations, and edges define the hierarchical relationships among them. Annotations are generated automatically through a pipeline that leverages pretrained multimodal large language models (MLLM) and object detection tools. The GBC structure enhances traditional image captions by providing both detailed descriptions and structural information, improving model performance in downstream tasks. All data is provided in JSON lines format, containing image URLs, bounding boxes, and caption annotations.
In TIPO, we only utilize the root node from GBC10M, since it provides both detail and short captions.
@@ A.3 Coyo HD 11M
The Coyo HD 11M dataset [12] consists of 11.4 million high-density, high-definition images, with 22.8 million synthetic captions generated from the Coyo-700M dataset. The “HD” refers to both the high resolution of images (filtered to have at least 512 pixels on the shortest edge) and the high concept density, addressing the issues found in previous alt-text image pair datasets, which often included low-quality or low-concept images. Synthetic captions for this dataset were produced using the LLaVA-Next-8B model [37] built upon LLaMA 3 [3] with post-processing for conciseness and clarity. This dataset provides high-quality visual and textual pairs, ideal for training advanced vision-language models.
In TIPO, we utilize the short/long captions, booru tags, and open images tags from Coyo HD 11M.
@@ A.4 Stable Diffusion XL
The Stable Diffusion XL [48] improves upon earlier versions [54] with a significantly larger UNet backbone and dual text encoders, CLIP ViT-L [49] and OpenCLIP ViT-bigG [30], to enhance text conditioning. It supports resolutions up to 1024×1024 and is designed for diverse image generation. Although explicit dataset details are not provided, figures in the paper confirm that SDXL accepts both natural language (NL) prompts and tags as input.
A.5 Stable Diffusion 3.5
The Stable Diffusion 3.5 (SD-3.5) incorporates a cutting-edge architecture, MMDiT [21], along with the Rectified Flow [38, 4, 35] formulation for improved text-to-image generation. The model utilizes triple text encoders—CLIP/ViT-L, OpenCLIP/ViT-G, and T5-XXL [50]—for robust text conditioning. SD-3.5 supports resolutions up to 1024×1024 and employs a 50/50 mix of original and CogVLM-generated captions for training. While explicit dataset details are not provided, figures in the paper confirm its capability to process both natural language prompts and tags.
In this study, we use the SD-3.5-Large [^8] https://huggingface.co/stabilityai/stable-diffusion-3.5-large (8B parameters) with FP8 inference on RTX 3090 or RTX 4090 GPUs.
@@ Appendix B Implementation Details
In this appendix, we provide all the necessary details including our dataset construction process, model configurations, inference pipeline and the model’s property which are not mentioned in Section 4.3 and 4.4.
@@ B.1 Training Data Construction
To systematically control output prompt length, we implement a structured length categorization system using special length tags. These tags enforce specific constraints on both tag counts and natural language sentence lengths. For instance, the <long> tag specifies that the corresponding prompt must contain between 36 and 52 tags (inclusive), accompanied by 4 to 8 sentences of natural language description. We define four distinct length categories, each with strict bounds for both tag count and sentence length.
@@
    This approach encourages the model to handle varying metadata positions and availability, while maintaining the ability to infer metadata relationships from the content descriptions.
@@
  • Natural Language: For natural language descriptions exceeding length limitations, we employ selective sentence removal, targeting middle sentences to preserve both context-setting opening sentences and concluding details. This maintains coherent narrative flow while meeting target length requirements.
@@ B.2 Training Settings and Model Configurations
Table 5: Training settings and details for TIPO-100M, 200M, and 500M. The datasets include CoyoHD11M (Coyo), GBC10M (GBC), and Danbooru2023 (Dan). Stage 2 additionally incorporates Pixtral [1] to generate NL captions from Danboodu2023 dataset.
@@
TIPO models are trained in multiple stages. Table 5 summarizes the configurations for pretraining and fine-tuning TIPO-100M, TIPO-200M, and TIPO-500M. Both pretraining and finetuning was conducted on datasets like Danbooru2023 [45], GBC10M [29], and CoyoHD11M [12].
@@ B.3 Inference Pipeline
The TIPO inference pipeline is designed to handle various input types and scenarios, combining different tasks to achieve refinement or expansion of both tag-based prompts and natural language prompts. Figure 8 illustrates this comprehensive workflow.
Our framework processes tags and natural language inputs separately, allowing for specialized handling of each input type. This flexible pipeline allows TIPO to adapt to various input scenarios, whether the user provides tags, natural language descriptions, or both. By leveraging different task combinations, TIPO ensures that both tag-based and natural language prompts are optimized, resulting in more detailed and effective inputs for text-to-image models.
@@ B.4 Model Speed Comparison
Table 6: Model performance comparison across different hardware platforms. Tokens per second (tok/sec) represents the average generation speed, while generation time (gen time) shows the average time required for a complete two-step prompt optimization process.
@@ Appendix C Evaluation Statistics
C.1 In-domain test regard to scenery tag
@@
The box plot and Kernel Density Estimation (KDE) plot shown in Figure 9 presents the aesthetic scores and AI corruption scores for results from the scenery tag test described in Section 5.2. The analysis clearly demonstrates that TIPO significantly outperforms all other methods, with a substantial margin of improvement.
@@ C.2 In-domain prompt generation test
Figures 10 and 11 display the box plots and KDE plots of aesthetic scores and AI corruption scores obtained from the In-domain prompt generation test detailed in Section E.2. While the box plots reveal that the differences in performance between various methods are subtle, the AI corruption scores provide valuable insights. Specifically, these scores indicate that implementations supported by TIPO produce more stable output images compared to other methods.
C.3 Out of focus evaluation
@@
Figures 12 and 13 present similarity matrices for different prompt generation methods and their corresponding aesthetic outputs on SD3.5-Large  [22]. A matrix with predominantly lower similarity values (overall brighter appearance) indicates high diversity across all generated images, while higher similarity values (overall darker appearance) suggest consistent but less diverse outputs.
For quantitative diversity metrics, please refer to Table 3 in Section 5.
@@ Appendix D TIPO example
In this appendix, we provide some text example of TIPO’s input and output.
TIPO Format template User Input:
1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, masterpiece, absurdres, newest, safe A girl sits in a cozy cafe, cradling a cup of coffee in her hand
Formatted TIPO Input For Expand Tags:
meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|short_to_tag|> short: A girl sits in a cozy cafe, cradling a cup of coffee in her hand tag: 1girl, solo
Formatted TIPO Output after Expand Tags and Expand Natural Prompt:
meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|tag_to_long|> tag: 1girl, solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding long: A girl sits in a cozy cafe, cradling a cup of coffee in her hand. The cafe has large windows with green plants on the walls and a wooden table in front of her. Behind her is a staircase leading to another room. The overall atmosphere of the image is serene and inviting.
Formatted Output For Text-to-Image
1girl, ciloranko, maccha \(mochancc\), ningen mame, ask \(askzy\), solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding, A girl sits in a cozy cafe, cradling a cup of coffee in her hand. Behind her is a staircase leading to another room. The cafe has large windows with green plants on the walls and a wooden table in front of her. The overall atmosphere of the image is serene and inviting. masterpiece, newest, absurdres, safe
Refer to caption User Input Refer to caption TIPO Output
@@ E.2 In-domain prompt generation test
Appendix F Topic Distribution Visualization

### v3

@@ A.1 Danbooru2023
The Danbooru2023 dataset [69, 68, 46] is an extensive collection of images and their corresponding tags, compiled from the Danbooru image board. This dataset includes images annotated with particular and detailed tags, providing a rich resource for training both the Text-to-Image (T2I) and Large Language Models (LLMs) involved in the TIPO framework. The dataset contains data up to image ID 7,349,999, encompassing various visual content with granular annotations. These annotations allow for creating nuanced and precise prompts, ensuring that longer, more detailed prompts can indicate subsets of shorter prompts.
@@
  • Rich Annotations: Detailed tags differentiate subtle variations, crucial for specific image generation.
  • Large Volume: Extensive dataset size ensures diverse training examples.
  • Tag-Based Prompting: Refined prompts from detailed tags enhance image generation accuracy.
@@ A.2 GBC10M
The GBC10M dataset [29] is a large-scale collection of 10 million images sourced from CC12M [13], annotated using the Graph-Based Captioning (GBC) approach. Each image is represented by a graph where nodes correspond to object regions, compositions, and relations, and edges define their hierarchical relationships. Annotations are generated automatically through a pipeline leveraging pretrained multimodal large language models (MLLM) and object detection tools. The GBC structure enhances traditional image captions by providing detailed descriptions and structural information. Data is provided in JSON lines format, including image URLs, bounding boxes, and captions.
In TIPO, only the root node captions from GBC10M are utilized for concise yet descriptive prompts.
@@ A.3 Coyo HD 11M
The Coyo HD 11M dataset [12] consists of 11.4 million high-resolution, high-concept-density images paired with 22.8 million synthetic captions generated from the Coyo-700M dataset. Images maintain a minimum of 512 pixels on the shortest edge to ensure high visual quality. Captions, generated with the LLaVA-Next-8B model [38] based on LLaMA 3 [3], undergo post-processing for conciseness and clarity.
TIPO uses short and long captions, booru tags, and open image tags from this dataset.
@@ A.4 Stable Diffusion XL
Stable Diffusion XL (SDXL) [50] improves upon earlier models [56] with a more considerable UNet backbone and dual text encoders (CLIP ViT-L [51] and OpenCLIP ViT-bigG [31]), enhancing text conditioning. Supporting resolutions up to 1024×1024, SDXL accepts natural language prompts and tags, suitable for diverse image generation.
A.5 Illustrious
Illustrious is a series of fine-tuned Stable Diffusion XL models primarily trained on the Danbooru2023 dataset. In this study, we specifically employ the v3.5 version variant with v-parameterization [58], which is notable for its extensive incorporation of natural language prompts. The inclusion of both tag-based and natural language formats allows Illustrious to leverage a broad range of semantic knowledge for image generation.
Within TIPO, we perform an ablation study to analyze the effectiveness of different prompting strategies—namely extended tags versus natural language prompts—to identify which approach contributes most significantly to enhanced image generation performance.
A.6 Stable Diffusion 3.5
Stable Diffusion 3.5 (SD-3.5) incorporates the MMDiT architecture [20] and the Rectified Flow formulation [39, 4, 36] for improved text-to-image generation. Utilizing triple text encoders (CLIP/ViT-L, OpenCLIP/ViT-G, T5-XXL [52]), SD-3.5 supports resolutions up to 1024×1024 and uses a 50/50 mix of original and CogVLM-generated captions. Figures confirm the capability to process both natural language prompts and tags.
This study employs SD-3.5-Large [^6] https://huggingface.co/stabilityai/stable-diffusion-3.5-large (8B parameters) with FP8 inference on RTX 3090 or RTX 4090 GPUs.
@@ Appendix B Implementation Details
In this appendix, we provide all the necessary details including our dataset construction process, model configurations, inference pipeline and the model’s properties not mentioned in Section 4.2 and 4.3.
@@ B.1 Training Data Construction
To systematically control output prompt length, we implement a structured length categorization system using unique length tags. These tags enforce specific constraints on tag counts and natural language sentence lengths. For instance, the <long> tag specifies that the corresponding prompt must contain between 36 and 52 tags (inclusive), accompanied by 4 to 8 sentences of natural language description. We define four distinct length categories, each with strict bounds for tag count and sentence length.
@@
    This approach encourages the model to handle varying metadata positions and availability, while maintaining the ability to infer metadata relationships from content descriptions.
@@
  • Natural Language: For natural language descriptions exceeding length limitations, we employ selective sentence removal, targeting middle sentences to preserve context-setting opening sentences and concluding details. This maintains coherent narrative flow while meeting target length requirements.
@@ B.2 Training Settings and Model Configurations
Table 6: Training settings for TIPO models. The datasets include CoyoHD11M (Coyo), GBC10M (GBC), and Danbooru2023 (Dan). Stage 2 additionally incorporates Pixtral [1] to generate NL captions from Danboodu2023 dataset.
@@
TIPO models are trained in multiple stages. Table 6 summarizes the configurations for pretraining and fine-tuning TIPO-100M, TIPO-200M, and TIPO-500M. Both pretraining and fine-tuning was conducted on datasets like Danbooru2023 [46], GBC10M [29], and CoyoHD11M [12].
@@ B.3 Inference Pipeline
The TIPO inference pipeline is designed to handle various input types and scenarios, combining different tasks to refine or expand both tag-based and natural language prompts. Figure 9 illustrates this comprehensive workflow.
Our framework processes tags and natural language inputs separately, allowing for specialized handling of each input type. This flexible pipeline allows TIPO to adapt to various input scenarios, whether the user provides tags, natural language descriptions, or both. By leveraging different task combinations, TIPO ensures that tag-based and natural language prompts are optimized, resulting in more detailed and effective input for text-to-image models.
@@ B.4 Model Speed Comparison
Table 7: Model performance comparison across different hardware platforms. Tokens per second (tok/sec) represents the average generation speed, while generation time (gen time) shows the average time in seconds required for a complete two-step prompt optimization process.
@@ Appendix C Evaluation Statistics
C.1 In-domain test regarding scenery tag
@@
The box plot and Kernel Density Estimation (KDE) plot displayed in Figure 10 illustrate the aesthetic scores and AI corruption scores from the scenery tag test described in Section 5.2. The analysis shows that TIPO significantly outperforms all other methods, demonstrating a considerable margin of improvement.
@@ C.2 In-domain prompt generation test
Figures 11 and 12 display the box plots and KDE plots of aesthetic scores and AI corruption scores obtained from the In-domain prompt generation test detailed in Section E.2. While the box plots reveal subtle differences in performance between various methods, the AI corruption scores provide valuable insights. Specifically, these scores indicate that implementations supported by TIPO produce more stable output images than other methods.
C.3 Out-of-domain evaluation
(a) The box plot for the Aesthetic Score result of out-of-focus test.
(b) The box plot for the AI Corrupt Score result of out-of-focus test.
(c) The KDE plot for the Aesthetic Score result of out-of-focus test.
(d) The KDE plot for the AI Corrupt Score result of out-of-focus test.
Figure 13: The distribution of aesthetic and AI corrupt score for out-of-focus test.
@@ C.3 Out-of-domain evaluation
Figures 14 and 15 present similarity matrices for different prompt generation methods and their corresponding aesthetic outputs on SD3.5-Large  [21]. A matrix with predominantly lower similarity values (brighter appearance) indicates high diversity among generated images, while higher values (darker appearance) suggest consistent but less diverse outputs. Please refer to Table 3 in Section 5.
C.4 Ablation Test
(a) The box plot for the Aesthetic Score result of NL ablation test.
(b) The box plot for the AI Corrupt Score result of NL ablation test.
(c) The KDE plot for the Aesthetic Score result of NL ablation test.
(d) The KDE plot for the AI Corrupt Score result of NL ablation test.
Figure 16: The distribution of aesthetic and AI corrupt score for NL ablation test.
(a) The box plot for the Aesthetic Score result of tags ablation test.
(b) The box plot for the AI Corrupt Score result of tags ablation test.
(c) The KDE plot for the Aesthetic Score result of tags ablation test.
(d) The KDE plot for the AI Corrupt Score result of tags ablation test.
Figure 17: The distribution of aesthetic and AI corrupt score for tags ablation test.
Figures 17 present the tag ablation test in the TIPO effect on the aesthetic score and AI Corrupt Score among the original tag, tag-extend and the tags TIPO. The box plot reveals that the tag TIPO is better than the original tag and the tag extend is the best. In detail, KDE plot reveals that the tag TIPO has a similar performance compared with the tag extend. Both of them are better than the original tag, which indicates that the tag TIPO aspect helps control corruption and promotes the aesthetic score.
@@ Appendix D TIPO example
In this section, we provide some text example of TIPO’s input and output.
TIPO Format Template User Input: 1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, masterpiece, absurdres, newest, safe A girl sits in a cozy cafe, cradling a cup of coffee in her hand Formatted TIPO Input for Expand Tags: meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|short_to_tag|> short: A girl sits in a cozy cafe, cradling a cup of coffee in her hand tag: 1girl, solo Formatted TIPO Output after Expand Tags and Expand Natural Prompt: meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|tag_to_long|> tag: 1girl, solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding long: A girl sits in a cozy cafe, cradling a cup of coffee in her hand. The cafe has large windows with green plants on the walls and a wooden table in front of her. Behind her is a staircase leading to another room. The overall atmosphere of the image is serene and inviting. Formatted Output for Text-to-Image: 1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding, A girl sits in a cozy cafe, cradling a cup of coffee in her hand. Behind her is a staircase leading to another room. The cafe has large windows with green plants on the walls and a wooden table in front of her. The overall atmosphere of the image is serene and inviting. masterpiece, newest, absurdres, safe Refer to caption User Input Refer to caption TIPO Output
Figure 18: An example of formatted content used for training and inference in TIPO.
@@ E.2 In-domain prompt generation test
Appendix F Human Preference
(a) ELO rating on Illustrious
(b) Win rate matrix on Illustrious
(c) ELO rating on SD3.5-medium
(d) Win rate matrix on SD3.5-medium
Figure 22: ELO ratings and win rate matrices across different experimental settings comparing five prompting methods (TIPO, Promptist, Promptextend, MagicPrompt, and Original) on three evaluation dimensions
(a) Full Win-Tie-Lose plot
(b) Illustrious Win-Tie-Lose plot
(c) SD3.5-medium Win-Tie-Lose
Figure 23: Win-Tie-Lose comparison across different experimental settings showing the relative performance of five prompting methods on prompt adherence, image quality, and aesthetic appeal metrics
(a) The UI of survey system before submitting the choices.
(b) The UI of survey system after submitting the choices.
Figure 24: Survey interface for human evaluation of image pairs, showing the evaluation process before submission (a) where users compare two images based on four metrics, and after submission (b) where the generated prompts for each image are revealed
Comparison Win Ratio (A:B) Proportion A $p$-value Significant?
Original vs. PromptExtend 45:66 0.405 0.0572 Marginally
Original vs. Tipo 28:80 0.259 $<0.0001$ Yes***
Original vs. Promptist 35:64 0.354 0.0046 Yes**
PromptExtend vs. Promptist 41:80 0.339 0.0005 Yes***
Promptist vs. Tipo 30:101 0.229 $<0.0001$ Yes***
PromptExtend vs. Tipo 30:90 0.250 $<0.0001$ Yes***
MagicPrompt vs. Promptist 14:34 0.292 0.0055 Yes**
MagicPrompt vs. Tipo 11:48 0.186 $<0.0001$ Yes***
MagicPrompt vs. Original 15:28 0.349 0.0660 No
MagicPrompt vs. PromptExtend 19:35 0.352 0.0402 Yes*
Table 8: Pairwise win rates and statistical significance (Overall Dimension). Significance levels: * $p<0.05$, ** $p<0.01$, *** $p<0.001$
We conducted a series of A/B tests to compare five prompt transformations, (TIPO, Promptist, Promptext, MagicPrompt, and Original(unmodified)), for two models, Illustrious, SD3.5-medium, which is known for both core word/natural language understanding. In total, we collected responses for $\sim$1,500 pairwise comparisons, from more than 20 anonymous evaluators. Each evaluation asked participants to compare two generated images—labeled A and B—and select which they preferred (or a tie) according to specific criteria (e.g., prompt adherence, image quality, or aesthetic appeal).
F.1 User Interface for Human Preference Evaluation
We developed a specialized survey interface to facilitate efficient and unbiased human evaluation of generated images. As illustrated in Figure 24, the interface presents evaluators with an original prompt and two corresponding images (labeled A and B) generated using different prompting methods. Before submission, users can see the original prompt in the center panel while the processed prompts used to generate each image remain hidden to prevent bias.
The evaluation framework requires participants to compare the image pairs across four distinct metrics: prompt adherence (how well the image follows the original prompt), image quality (detail and correctness), aesthetic appeal (color, composition, and style), and overall personal preference. For each metric, users can select one of three options: “A is better,” “A and B are equal,” or “B is better.”
When evaluators encounter image pairs that appear to be from different prompts or settings, they are instructed to click “Refresh” to obtain a new comparison. After submitting their evaluations, the interface reveals the transformed prompts used to generate each image, providing transparency about how the original prompt was modified by each method.
F.2 Extended Human Evaluation.
Participants assessed each image’s performance on prompt adherence, image quality, and aesthetic appeal, with visually shown unmodified and image pairs. TIPO exhibited superior outcomes in all comparison settings. Notably, it attained a 64.4% peak win rate (against MagicPrompt) under the Full scenario and 57.5% (also against MagicPrompt) under SD35-medium, emphasizing TIPO’s proficiency in generating images that closely follow prompt specifications while maintaining visual coherence.
F.3 ELO Ratings.
We computed theoretical ELO ratings from the aggregated pairwise comparisons to quantify overall performance differences among the five methods. The rating update rules were based on each pair’s binary outcome (win or lose), ignoring tie cases. The result is depicted in Figure 22, TIPO has secured the highest ELO rating over other models.
F.4 Human Preference ELO Method
We computed theoretical ELO ratings from human-judged pairwise preference data to quantitatively evaluate the relative performance of each prompting method. The ELO rating system, initially designed for ranking chess players, aggregates binary outcomes into numerical ratings representing comparative performance.
Pairwise Outcomes.
Human evaluators assessed comparisons between methods, resulting in one of three outcomes:
  • Method $i$ wins: assigned a score of $1$ for method $i$, and 00 for method $j$.
  • Method $j$ wins: assigned score $1$ for method $j$, and 00 for method $i$.
  • Tie: assigned score $0.5$ to both methods.
Conversion to ELO Differences.
Win and tie rates were converted to ELO rating differences using:
Adjusted Win Rate=Win Rate+Tie $\text{Adjusted Win Rate}=\text{Win Rate}+\frac{\text{Tie Rate}}{2}$
ELO $\text{ELO Difference}=400\times\log_{10}\left(\frac{\text{Adjusted Win Rate}}{1-\text{Adjusted Win Rate}}\right)$
To ensure numerical stability, extreme adjusted win rates were constrained as follows:
ELO $\text{ELO Difference}=\begin{cases}-800,&\text{Adjusted Win Rate}\leq 0.001\\ +800,&\text{Adjusted Win Rate}\geq 0.999\\ \end{cases}$
Calculating Method ELO Ratings.
Final ELO ratings were determined by averaging each method’s pairwise ELO differences and centering these averages around a baseline rating (e.g., $1000$):
$\text{ELO}_{\text{method}_{i}}=\text{Base Rating}+\left(\text{Average ELO Difference for method }i-\text{Overall Mean ELO Difference}\right)$
Interpretation of ELO Scores.
Methods with higher ELO scores consistently outperform lower-scored methods. A rating difference of 400 points corresponds to a 90% expected win probability for the superior method.
F.5 Statistical Significance
As summarized in Table 8, we conducted two-sided binomial and McNemar’s tests (p <0.05) to assess the statistical significance of observed differences. The result confirms that TIPO’s advantages are unlikely to be explained by random variation, which also supports a consistent performance hierarchy: TIPO ranks highest, followed by Promptist, PromptExtend, Original, and MagicPrompt. Collectively, these findings illustrate TIPO’s robust, model-agnostic effectiveness and underscore the model-sensitivity of alternative methods, particularly Promptext and MagicPrompt.
F.6 Survey Response Examples
In this section we provided some responses of our human preference survey as reference.
Figure 25: Some survey responses on illustrious-3.5-vpred generated image with different prompt optimization method
Figure 26: Some survey responses on illustrious-3.5-vpred generated image with different prompt optimization method
Figure 27: Some survey responses on SD3.5-medium generated image with different prompt optimization method
Figure 28: Some survey responses on SD3.5-medium generated image with different prompt optimization method
F.7 Conclusion
The extended evaluations presented here reinforce TIPO’s standing as a reliable and effective prompt-optimization strategy. Its consistent performance gains under diverse model conditions highlight its potential for broad applicability, with its strong alignment with user-specified prompts, high image quality, and favorable aesthetic outcomes.
Appendix G Ablation Study on TIPO
In this section, we investigate the effect of incorporating TIPO (Tags + Inferred Prompt Objects) across various generation settings. Our primary goal is to validate whether additional structured information (e.g., core tags and minimal spatial/contextual cues) can improve image quality, reduce artifacts.
G.1 Experimental Setup
Prompt Variants.
To systematically analyze TIPO’s contribution, we consider four types of input prompts improvement task:
  1. 1.
    Tag $\to$ More core words: Given an initial set of core words, generate more refined or expanded core words.
  2. 2.
    NL $\to$ More NL: Given a short natural language (NL) description, elaborate into a richer NL prompt.
  3. 3.
    Tag $\to$ (More core words + NL): Combine expanded tags with a corresponding NL description derived from them.
  4. 4.
    NL $\to$ (More NL + core words): Use the NL prompt to add relevant tags, forming a mixed prompt of NL plus core words.
In each case, we compare the baseline prompts (without TIPO cues) against prompts incorporating TIPO’s structured, tag-based critical information and minimal spatial hints.
Data Preparation.
We start by randomly sampling core words from a word table to represent a diverse range of topics (e.g., objects, environments, descriptors). Additionally, for each word set, we generate a corresponding short NL sentence using a compact language model (GPT4o-mini). Overall, the six prompt variants are tested on 4,000 images, ensuring a balanced comparison.
Inference Procedure.
Prompts are fed into our image-generation pipeline under identical model settings (classifier free guidance, sampler, steps, etc.), using the v-parameterized variant of Illustrious v3.5[49]. We focus on how TIPO modifications alter the generation outcomes and whether they introduce additional computational overhead.
G.2 Evaluation Metrics
(a) Corrupt Score Distributions
(b) Aesthetic Score Distributions
Figure 29: Side-by-side comparison of Corrupt (left) and Aesthetic (right) score distributions across prompt types.
Aesthetic Score.
We employ an off-the-shelf aesthetic predictor to estimate image quality. In the following paragraph, we discuss the model’s bias.
AI Corruption Score.
Using an automated ’ AI corruption ’ detection model, we measure generation artifacts, such as distorted objects and unnatural shapes. Higher scores imply cleaner, more coherent outputs.
G.3 Results & Discussion
Impact on Aesthetics.
Figure 29(b) shows that TIPO-enhanced prompts generally achieve higher aesthetic scores than their non-TIPO counterparts, albeit with some variance. Notably, we observe a correlation between wider color ranges and higher aesthetic scores discussed in Figure 30, suggesting a bias toward more colorful or varied compositions.
Improvements in Corruption Score.
As shown in Figure 29(a), TIPO-based prompts yield significantly lower corruption scores, indicating fewer artifacts. We hypothesize that the additional spatial and contextual details encoded via TIPO help the model place objects more consistently.
(a) Saturation vs. Aesthetic Score
(b) Binned Saturation vs. Aesthetic
(c) Saturation vs. Corrupt Score
Figure 30: Scatter plots (left and right) and binned analysis (middle) showing the relationship between saturation and image metrics. We find a moderate positive correlation between saturation and aesthetic score (Pearson $r=0.2821$), particularly at lower saturation ranges, based on 24k samples. However, saturation shows no notable correlation with corrupt score (Pearson $r=0.0125$).
G.4 Speed Test and Overhead Analysis
A key concern for production pipelines is whether TIPO generation imposes a substantial time overhead. We benchmarked prompt-generation inference on four smaller models, excluding any large proprietary LLMs. As illustrated in Table 9, the additional TIPO-related computation remains well below the image-generation time. Hence, even in a synchronous pipeline, TIPO prompt expansion does not constitute a bottleneck.
Table 9: Speed Test Results for TIPO and Other Prompt Methods
Method Model/Config # Runs Avg. Time (s) Std. Dev. (s)
TIPO LLaMA-500M 500 1.4207 1.0730
LLaMA-200M 200 1.0306 0.8982
LLaMA-100M 200 1.0078 0.9394
PROMPTIST GPT2-125M 1000 1.4593 0.2857
PROMPTEXTEND GPT2-125M 1000 1.3849 0.2151
MAGICPROMPT GPT2-125M 1000 1.1398 0.4043
Memory Footprint.
We also confirm that TIPO’s overhead in terms of VRAM usage is minimal. The practical adoption of TIPO in pipelines has shown no critical memory concerns, which aligns with our measurements.
G.5 Conclusion of Ablation
Our experiments suggest that TIPO (1) consistently lowers AI corruption artifacts, (2) can boost aesthetic scores, and (3) remains computationally inexpensive. The improvements in metrics support the viability of TIPO prompts for real-world image-generation tasks. In short, a concise natural language prompt with core tag-based critical information appears to be an effective, suggested form for most use cases.
Appendix H Topic Distribution Visualization